ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 119, 2000

Ballico, E. Trigonal Gorenstein curves and special linear systems  143

Bar-Natan, D., S. Garoufalidis, L. Rozansky and D. P. Thurston Wheels, wheeling, and the Kontsevich integral of the unknot

217

Boyland, P. New dynamical invariants on hyperbolic manifolds  253

Džamonja and S. Shelah Erratum: “♣ does not imply the existence of a Suslin tree”, Israel Journal of Mathematics, Vol. 113, 1999, pp. 163–204

379

Garoufalidis, S. See D. Bar-Natan

Gustafsson, B. and M. Putinar Linear analysis of quadrature domains. II

187

Knaf, H. Divisors on varieties over valuation domains  349

Lima, Å., O. Nygaard and E. Oja Isometric factorization of weakly compact operators and the approximation property

325

Martinez, P. Precise decay rate estimates for time-dependent dissipative systems

291

Masters, J. D. Length multiplicities of hyperbolic 3-manifolds  9

Nygaard, O. See Å. Lima

Oja, E. See Å. Lima

Polterovich, I. Heat invariants of Riemannian manifolds  239

Putinar, M. See B. Gustafsson

Rozansky, L. See D. Bar-Natan

Shelah, S. See M. Džamonja

Solan, E. Stochastic games with two non-absorbing states  29
Sonn, J. Free pro-p groups as Galois groups over ℚ(p)(t)  1

Thurston, D. P. See D. Bar-Natan

Vieille, N. Small perturbations and stochastic games  127
Vieille, N. Two-player stochastic games I: A reduction  55

Vieille, N. Two-player stochastic games II: The case of recursive games

93

Zhang, G. Invariant differential operators on Hermitian symmetric spaces and their eigenvalues 157